How to changed mixed fractions to their simplest form??
3 10/15
5 12/18

Answer by Alan3354(69443)

3 10/15 = 45/15 + 10/15 = 55/15
= 11/3
---------------------------------
5 12/18 = 5 + 2/3
= 15/3 + 2/3
= 17/3